OpenStudy (anonymous):

Find the volume of the solid generated by rotating the area bounded by the curves y=cotx and y=2cosx for 0<=x<=Pi/2, around the line y=-1

OpenStudy (anonymous):

i come to this equation \[\int\limits_{0}^{\Pi/2} \Pi(2cosx+1)^2-\Pi(Cotx+1)^2\]

OpenStudy (anonymous):

should this be right?

sam (.sam.):

Is that capital pi or small pi?

OpenStudy (turingtest):

you're gonna need two integrals I believe

OpenStudy (anonymous):

its just pi

OpenStudy (turingtest):

but... x cannot be zero another strange problem from your prof

OpenStudy (anonymous):

haha really!? again?

OpenStudy (anonymous):

thats why i post this to see if im doing it right

OpenStudy (anonymous):

i try to do it on wolfram but its not letting me

OpenStudy (turingtest):

cot is undefined at x=0 so the volume is infinity unless this is some kind of improper integral problem I haven't figure out yet

OpenStudy (turingtest):

look at the area between the graphs http://www.wolframalpha.com/input/?i=plot+y%3Dcotx%2Cy%3D2cosx

OpenStudy (turingtest):

if it was from pi/6 to pi/2 that would be ok...

OpenStudy (turingtest):

maybe it's a right-hand improper integral, let me try that...

OpenStudy (anonymous):

seems like the two function doesnt intersect at 0

OpenStudy (turingtest):

yes I was just coming to that conclusion from my own angle http://www.wolframalpha.com/input/?i=integral+from+0+to+pi%2F6+cot%5E2x

OpenStudy (turingtest):

(^this integral is in the answer i would figure)

OpenStudy (turingtest):

and a right-hand limit is not helping me, because the integral of cot^2 is -x-cotx so cotx is still in the problem :/

OpenStudy (anonymous):

yea thats was my problems too

OpenStudy (anonymous):

so in order to do this problem the limits have to be change?

OpenStudy (turingtest):

yeah, or it needs a different boundary unless I'm missing something major you have a very difficult problem, as even the CPV in this problem is infinity according to wolfram

OpenStudy (anonymous):

ok another question to ask the prof about. thanks turing?

OpenStudy (turingtest):

welcome, but feel free to get a second opinion

OpenStudy (anonymous):

i did the whole problem but the answer.. im not confident with it

OpenStudy (turingtest):

how did you integrate cot ?

OpenStudy (turingtest):

btw, the way I see it the integral is this\[V=\pi\int_{0}^{\frac\pi6}\cot^2x-4\cos^2xdx+\pi\int_{\frac\pi6}^{\frac\pi2}4\cos^2x-\cot^2xdx\]

OpenStudy (turingtest):

that's pi/6 in the middle there

OpenStudy (dumbcow):

yeah change your limits to pi/6 to pi/2 because that is where the functions intersect creating the region to be revolved around y=-1 i think the 0<x<pi/2 , is to restrict the domain of the intersecting points

OpenStudy (turingtest):

if that is the case dumbcow the problem is very ill-written, you must agree and it is not just 0<x<pi/2 but\[0\le x\le\frac\pi2\]if you want the region between the graphs in that interval you are in trouble.

OpenStudy (anonymous):

hey turing just woke up and started working on this problem again.. so i will have to change the limits to Pi/6 and Pi/2

OpenStudy (anonymous):

can you tell me how did you get the equation to intergrate?

OpenStudy (anonymous):

i thought its outside radius^2 - inside radius^2

OpenStudy (turingtest):

yeah I messed up and forgot the 1, so from pi/6 top pi/2 it should be\[\large V=\pi\int_{\frac\pi6}^{\frac\pi2}(2\cos x+1)^2-(\cot x+1)^2dx\]which I think is what you had

OpenStudy (anonymous):

yea thats what i got hehe

OpenStudy (anonymous):

and just integrate on right?

OpenStudy (turingtest):

yep, nothing problematic-looking.

OpenStudy (turingtest):

looks like you might save a little trouble by using the formula cot^2+1=csc^2 on the after you expand the second term

OpenStudy (turingtest):

on the right*

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

time to eat breakfast!

OpenStudy (turingtest):

enjoy :D